OpenStudy (anonymous):

A charge of 1.82E-9 C is placed at the origin, and a charge of 4.32E-9 C is placed at x = 1.26 m. Find the position at which a third charge of 3.05E-9 C can be placed so that the net electrostatic force on it is zero. I am assuming that the charges are in a straight line. There was not an illustration with it.

OpenStudy (anonymous):

yes charges are in a straight line & the third charge placed between to charges & near the smaller charge ,do you understand?

OpenStudy (turingtest):

assuming all charges are positive\[\sum F=k\frac{q_1q_2}{r_{12}^2}-k\frac{q_2q_3}{r_{23}^2}=0\]solve for r23